OpenStudy (anonymous):

PLEASE HELP ME!! Diane can choose Plan A or Plan B for her long distance charges. For each plan, cost (in dollars) depends on minutes used (per month) as shown below.

OpenStudy (danjs):

Gonna be the same type of problem involving a lines. COST PLAN A = k*(Minutes used) COST PLAN B = j*(Minutes Used) k and j are the slopes of each of the lines for plan A and B

OpenStudy (anonymous):

PLAN A Here r the points (0,0) (20,2) (40,4) (60,6) (80,8) (100,10) (120,12) (140,14) (160,16) (180,18) (200,20) (220,22) (240,24) (260,26)

OpenStudy (anonymous):

PLAN B Here r the points: (0,12) (20,12) (40,12) (60,12) (80,13) (100,14) (120,15) (140,15) (160,16) (180,17) (200,18) (220,19) (240,19) (260,20)

OpenStudy (anonymous):

wait wait

OpenStudy (danjs):

sorry i have to leave, i will be back later on ...

OpenStudy (anonymous):

PART A: If Diane makes 60 minutes of long distance calls for the month, which plan costs more? PLAN A or PLAN B How much more does it cost than the other plan? PART B: for what number of long distance minutes do the two plans cost the same? if the same spent on long distance calls is more than this amount, which plan costs more? PLAN A OR PLAN B?

OpenStudy (anonymous):

@Michele_Laino please help me?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@pooja195 please help me

OpenStudy (anonymous):

@DanJS help me please this is my last question =(